Compute for the maximum height to be reached by the cannon


ball.

1
Expert's answer
2022-03-03T13:23:06-0500

The maximum height depends on the initial speed (see https://courses.lumenlearning.com/boundless-physics/chapter/projectile-motion/):


"h = \dfrac{u_0^2\sin^2\theta}{2g}"

where "u_0" is the initial speed, "\theta" is the launch angle, "g" is the gravitational acceleration.